APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday June 2, 2009 (MJD 54985) ---=== OBSERVING TEAM ===--- Viktor Malanushenko Olena Malanushenko ---=== OBSERVING PLAN ===--- Science! 2 Jun 2009 54985.0 0.71 MARVELS 7.5 hours ---=== OBSERVING SUMMARY ===--- The weather was good at the begining of observations but dramatically changed in a while with wind and humidity up, and temperature down. But still have chance to observe two plates: Cart 9 plate 3280B, HD89744, 50 min exposure Cart 1 plate 3277B, GJ436, 50 min exposure Missing fibers 89-90 on Cartridge 9, side B, blue bandle #6, blue, magenta ---=== OBSERVING LOG ===--- Afternoon: --------- All checked and OK. 01:12Z opened doors, louvers, start ventilation. Night: ----- 02:05Z Enclosure is off Cart 9 plate 3280B, Marvels FK5 Calibrations MARVELS1.2009-06-03T02:45:44.640.fits MARVELS1.2009-06-03T02:47:55.770.fits MARVELS1.2009-06-03T02:52:31.790.fits Stars, science, HD89744, 50 min exposure MARVELS1.2009-06-03T03:01:05.640.fits =32.72, seeing 1.2" Missing fibers 89-90. Changing cartridge Calibration MARVELS1.2009-06-03T03:53:23.850.fits MARVELS1.2009-06-03T03:55:31.880.fits MARVELS1.2009-06-03T04:00:08.490.fits Cart 1 plate 3277B, GJ436, 50 min exposure MARVELS1.2009-06-03T04:08:12.050.fits =14, seeing 1.6". Clouds during the observations. Changing the cartridge. Calibrations: MARVELS1.2009-06-03T05:01:30.980.fits MARVELS1.2009-06-03T05:03:40.430.fits MARVELS1.2009-06-03T05:08:14.900.fits Cart 5 plate 4314A, FK4. The weather is changed gramaticly for last 10-15 minuts. Wind is up to 38-40 mph, humidity up from 20% to 80%, and the different is around 3C. We did not know how worse it could be soon and closed. 06:50Z stopped ventilation and enclosure was on. Wait for weather improvment. logTool is hanged and information about focuses and offsets have been lost. Closed iop and opened again, and restored information manually. The wind went down but the humidiry did not drop during the night. 04:15 Z - endNight. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 61526, plate 3277 ------- 00:45 61526 10.0 flat 00:48 61527 2.0 arc 00:51 61528 3.0 arc 00:54 61529 3.0 arc ----- sequence 61530, plate 3495 ------- 00:58 61530 10.0 flat 01:01 61531 2.0 arc 01:04 61532 2.0 arc 01:06 61533 3.0 arc ----- sequence 61534, plate -9999 ------- 01:57 61534 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:14Z 60 20 14 277 (W) 1436 - 00:48Z 61 18 9 284 (WNW) 1448 - 01:20Z 61 19 7 253 (WSW) 1470 - 01:50Z 59 19 2 276 (W) 1483 - 02:03Z 59 19 0 250 (WSW) 1435 - 02:35Z 58 19 0 160 (SSE) 1441 - 03:06Z 59 18 1 261 (W) 1537 - 03:36Z 59 16 5 271 (W) 1386 - 05:12Z 48 39 31 72 (ENE) 2398 - 05:42Z 47 41 26 70 (ENE) 2112 - 06:12Z 45 41 28 68 (ENE) 1269 - 06:44Z 44 41 33 68 (ENE) 996 - 07:16Z 44 41 21 64 (ENE) 1117 - 07:46Z 43 41 38 73 (ENE) 1401 - 08:17Z 43 42 21 74 (ENE) 1352 - 08:49Z 44 41 22 76 (ENE) 1715 - 09:21Z 44 41 22 69 (ENE) 1699 - 09:51Z 44 40 20 72 (ENE) 1572 - 10:26Z 44 40 18 72 (ENE) 1868 - ---=== TELESCOPE STATUS ===--- 01:12Z - Started enclosure ventilation 02:05Z - Opened enclosure 06:50Z - Closed enclosure. Ventilation off. 10:40Z - Telescope at 30 deg alt. Cart 9 mounted Counterweights at 270 Spectro autofill connected No interlocks bypassed Spectro dewar weight 165 lb; 20 psi Imager dewar disconnected ---=== SOFTWARE USED ===--- IOP/SOP: v4_39_0 Watcher: v2_36_0 MCP: v6_6_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.10.0 2009-01-12 sdssProcedures: SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0240 -0.0280 Axial B 0.0300 0.0470 Axial C -0.0290 -0.0360 Trans D -0.7610 -0.7600 Lateral E 10.5791 10.5791 Lateral F 12.1793 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10471. -76. 5755. 8616. 10374. 10023. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: 15.5 Alt.: 29.701427 MIGS TONIGHT NOMINAL Axial A 18.5870 18.5840 Axial B 15.6330 15.6220 Axial C 13.8060 13.8150 Trans D 15.1020 15.1080 GALILS Commanded: 1626660. 1528833. 1595617. -2950. -4550. Actual: 1626627. 1528808. 1595583. -2945. -4561. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1257.02 -12.00 -28.00 -39.19 94.70 ---=== PROBLEMS IN DETAIL ===--- Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00061525.fit -5.9 -4.3 -2.6 2.0 7.3 9.4 10.6 sdR-b2-00061525.fit -8.0 -6.3 -4.2 1.1 7.3 9.7 11.0 sdR-r1-00061525.fit -6.4 -5.2 -3.4 2.4 7.9 9.6 12.2 sdR-r2-00061525.fit -7.5 -5.6 -4.3 1.8 8.1 9.3 12.3 --------------- sp1> Coll_motor_A 2300 Coll_motor_B 2299 Coll_motor_C 1698 Temp_Median 16.7 sp2> Coll_motor_A 1301 Coll_motor_B 1293 Coll_motor_C 900 Temp_Median 15.5 ----------------- 00:43Z postCalib 00:49Z goSpecFocus ----------------- bias 61525 - ok FF 61526 - ARC 61527 - ARC1 61528 to collimate ARC2 61529 to collimate wrong plate number: cart 1 pl 3277 instead cart 2 pl 3495 ! ------------------ 00:57Z postCalib Cart 2 plate 3495 FF 61530 - ok ARC 61531 - ok 01:02Z ARC1 61532 to collimate ARC2 61533 to collimate ----------------------- collimate, 61532 % Compiled module: COLLIMATE. Reading /data/spectro/54985/sdR-b1-00061532.fit Reading /data/spectro/54985/sdR-b1-00061533.fit COLLIMATE: Min offset = -0.22 pix COLLIMATE: Max offset = 0.89 pix COLLIMATE: RMS across CCD = 0.29 pix COLLIMATE: Mean offset = 0.27 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -9.6 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54985-b1-00061532.log PostScript file = Collimate-54985-b1-00061532.ps Reading /data/spectro/54985/sdR-b2-00061532.fit Reading /data/spectro/54985/sdR-b2-00061533.fit COLLIMATE: Min offset = 0.01 pix COLLIMATE: Max offset = 0.72 pix COLLIMATE: RMS across CCD = 0.21 pix COLLIMATE: Mean offset = 0.39 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -14.0 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54985-b2-00061532.log PostScript file = Collimate-54985-b2-00061532.ps Reading /data/spectro/54985/sdR-r1-00061532.fit Reading /data/spectro/54985/sdR-r1-00061533.fit COLLIMATE: Min offset = -0.81 pix COLLIMATE: Max offset = 0.49 pix COLLIMATE: RMS across CCD = 0.35 pix COLLIMATE: Mean offset = -0.24 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2217 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 2217" Log file = Collimate-54985-r1-00061532.log PostScript file = Collimate-54985-r1-00061532.ps Reading /data/spectro/54985/sdR-r2-00061532.fit Reading /data/spectro/54985/sdR-r2-00061533.fit COLLIMATE: Min offset = -1.25 pix COLLIMATE: Max offset = -0.02 pix COLLIMATE: RMS across CCD = 0.37 pix COLLIMATE: Mean offset = -0.43 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 3971 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 3971" Log file = Collimate-54985-r2-00061532.log PostScript file = Collimate-54985-r2-00061532.ps -----------